OpenStudy (anonymous):

What is the value of y when x = -1? 5x + 4y = 11

OpenStudy (callisto):

Put x=-1 into 5x + 4y = 11 5(-1) + 4y = 11 -5 +4y =11 add 5 to both sides 5+ (-5) + 4y = 11+5 4y = 16 Divide both sides by coefficient of y, what do you get from here?

OpenStudy (anonymous):

THANKS!